# Leveraging Knowledge Graph-Based Human-Like Memory Systems to Solve Partially Observable Markov Decision Processes

### v1

@@ Abstract
Humans observe only part of their environment at any moment but can still make complex, long-term decisions thanks to our long-term memory system. To test how an AI can learn and utilize its long-term memory system, we have developed a partially observable Markov decision processes (POMDP) environment, where the agent has to answer questions while navigating a maze. The environment is completely knowledge graph (KG) based, where the hidden states are dynamic KGs. A KG is both human- and machine-readable, making it easy to see what the agents remember and forget. We train and compare agents with different memory systems, to shed light on how human brains work when it comes to managing its own memory systems. By repurposing the given learning objective as learning a memory management policy, we were able to capture the most likely belief state, which is not only interpretable but also reusable. Our code is open-sourced at https://github.com/humemai/agent-room-env-v2-lstm.
@@ 1 Introduction
The cognitive science theory suggests that humans have a long-term memory system (Tulving 1983, 1985; Tulving and Thomson 1973). This allows humans to solve various tasks, although they can only observe the world partially. For example, when trying to navigate through a familiar city, humans rely on stored memories of the layout and landmarks. This memory system enables the recall of past experiences, aiding in decision-making and problem-solving even when immediate sensory input is limited. Similarly, when answering a question, humans draw on long-term memory to retrieve relevant facts, concepts, and previous knowledge.
Although there have been many studies, we still lack a comprehensive understanding of how our long-term memory system is managed. Research has identified mechanisms for encoding, storing, and retrieving memories, but the precise regulation and optimization of these processes remain unclear (Artuso and Palladino 2019; Squire 2009).
@@
  • Optimal Memory Management Policy: We show that the best agent is the one that repurposes the given learning objectives as learning a memory management policy, allowing itself to capture the most likely belief state, which is not only interpretable but also reusable.
@@ 2.1 A Machine With Human-Like Memory Systems
Recent advances in cognitive science and AI have inspired the development of machines with memory systems inspired by human memory. Kim et al. (Kim et al. 2022, 2023) introduced an agent model that explicitly incorporates both semantic and episodic memory systems, demonstrating that such an agent outperforms those with only one type of memory system. Their environment is designed to challenge the agent’s ability to encode, store, and retrieve memories effectively.
@@
To computationally model episodic and semantic memories, Kim et al. (Kim et al. 2022, 2023) add relation qualifier key-value pairs to RDF (Resource Description Framework) triples to make them RDF quadruples (Lanthaler, Wood, and Cyganiak 2014; Hartig et al. 2024). For example, given a memory triple (Agent, atLocation, Library), they would turn this into something like (Agent, atLocation, Library, {timestamp: 42}), and (Agent, atLocation, Library, {strength: 2}), to turn them into episodic and semantic memories, respectively.
@@ 2.2 Reinforcement Learning (RL) in Partially Observable Markov decision process (POMDP)
RL is useful when supervision of every action is not feasible, such as in memory management. This is often formalized as a Markov Decision Process (MDP) or a Partially Observable MDP (POMDP), which account for partial observability.
@@
The environment starts in an initial state distribution. At time step $t$, the agent maintains a belief state $\mathrm{b}_{t}$, which is a probability distribution over the state space $\mathcal{S}$. The agent then chooses an action $\mathrm{a}_{t}\in\mathcal{A}$ based on its current belief state $\mathrm{b}_{t}$ according to a policy $\pi:\mathcal{B}\rightarrow\mathbb{P}(\mathcal{A})$, where $\mathcal{B}$ denotes the belief space. After taking action $\mathrm{a}_{t}\sim\pi(\mathrm{a}_{t}\mid\mathrm{b}_{t})$, the agent observes an observation $\mathrm{o}_{t+1}\in\Omega$. Using the observation $\mathrm{o}_{t+1}$ and the action $\mathrm{a}_{t}$, the agent updates its belief state to $\mathrm{b}_{t+1}$.
@@
where $r_{t}=R(\mathrm{s}_{t},\mathrm{a}_{t})$, $\mathrm{a}_{t}\sim\pi(\mathrm{a}_{t}\mid b_{t})$, $\mathrm{s}_{t+1}\sim T(\mathrm{s}_{t},\mathrm{a}_{t})$, and $\mathrm{o}_{t+1}\sim O(\mathrm{o}_{t+1}\mid\mathrm{s}_{t+1},\mathrm{a}_{t})$. The belief state $b_{t+1}(\mathrm{s}_{t+1})$ is updated using the current belief state $\mathrm{b}_{t}$, the action $\mathrm{a}_{t}$, and the observation $\mathrm{o}_{t+1}$, along with the observation model $O$ and the transition model $T$. The update rule is given by:
$b_{t+1}(\mathrm{s}_{t+1})=\frac{O(\mathrm{o}_{t+1}\mid\mathrm{s}_{t+1},\mathrm{a}_{t})\sum_{\mathrm{s}_{t}\in\mathcal{S}}T(\mathrm{s}_{t+1}\mid\mathrm{s}_{t},\mathrm{a}_{t})b_{t}(\mathrm{s}_{t})}{\sum_{\mathrm{s}^{\prime}\in\mathcal{S}}O(\mathrm{o}_{t+1}\mid\mathrm{s}^{\prime},\mathrm{a}_{t})\sum_{\mathrm{s}_{t}\in\mathcal{S}}T(\mathrm{s}^{\prime}\mid\mathrm{s}_{t},\mathrm{a}_{t})b_{t}(\mathrm{s}_{t})}$ (2)
Updating the belief state can be computationally expensive due to several factors. The belief state represents a probability distribution over the entire state space, which can be vast. Updating it requires summing over all possible states and applying transition and observation probabilities, making the process highly intensive.
A simple way to mitigate this computational expense is to use the history of the past $\tau$ observations $\bm{h}^{(\tau)}_{t}=(o_{t-\tau+1},\ldots,o_{t})$ as the estimated state $\hat{s}_{t}$, with which we learn the policy $\pi(\mathrm{a}_{t}\mid\mathrm{\hat{s}}_{t})$. This alleviates the need to maintain or update the belief state $\mathrm{b}_{t}$.
@@ 3.1 The Rooms Environment
The Rooms Environment provides a simulation for evaluating AI agents with long-term memory systems. Unlike the previous work, this environment consists of multiple interconnected rooms, each containing various types of objects. Static objects, such as beds and tables, remain fixed in their locations and do not move. Independent objects, like Bob and Alice, move between rooms according to probabilistic rules that dictate their movements. Dependent objects, which cannot move by themselves, may be transported by independent objects with a certain probability that depends on the current state of the room. Finally, the agent represents the RL agent that navigates between rooms based on its exploration policy. This configuration allows for the study of how agents manage and utilize long-term memory to make decisions in a dynamic and partially observable environment.
In this environment, hidden states at time $t$, $s_{t}$, are represented as dynamic KGs that capture the entire rooms and the locations of all objects within them. Each room and its contents form a sub-graph of this overall KG. Since the agent can only be present in one room at a time, it can only observe the objects within that particular room. This partial observability requires the agent to utilize its long-term memory.
@@
Figure 1 shows an example hidden state $s_{t}$ and partial observation $o_{t}$. In this example, there are 32 rooms, i.e., Office room, Kitchen, …, and 25 objects, i.e., Coat, Towel, …, Agent. Therefore, the number of all hidden states is $|S|=N_{\text{rooms}}^{(N_{\text{objects}}-N_{\text{static objects}})}=32^{25-8}\approx 3.8\times 10^{25}$, since every object is located in one of $N_{\text{rooms}}$ and the static objects do not move. The number of possible sub-graphs, i.e., partial observation, that the agent can observe is $2^{(N_{\text{objects}}-N_{\text{static objects}})}\times N_{\text{rooms}}=2^{25-4}\approx 6.7\times 10^{7}$, since the agent can either see or do not see an object in the room. Although the environment might look simple, we have a huge state space, which can be too much for typical table-based RL methods.
@@
In this example, the partial observation (sub-graph of the hidden dynamic KG) is { (Closet, north, Nursery), (Closet, east, Wall), (Closet, south, Wall), (Closet, west, Sauna), (Agent, atLocation, Closet) }. Then the hidden state is a much bigger KG that contains this sub-graph.
@@
(a) An illustration of a hidden state $s_{t}$ (in white) and partial observation $o_{t}$ (in gray).
@@
(b) A hidden state $s_{t}$ (in white) and partial observation $o_{t}$ (in gray) represented as a KG.
Figure 1: An example hidden state $s_{t}$ and partial observation $o_{t}$ of the Rooms environment.
@@
The exploration policy, $\pi^{\text{explore}}(a^{\text{explore}}|o_{t})$, allows the agent to navigate through the room by taking one of five possible actions: North, East, South, West, or Stay. As the agent explores, the environment intermittently poses questions about the object locations. Consequently, the agent must not only learn an effective exploration policy, $\pi^{\text{explore}}$, but also develop a robust question-answering function, $answer=f^{qa}(o_{t})$. The question-answering function also serves as the reward function: the agent receives a reward of +1 for each correctly answered question and 0 otherwise. Although straightforward, using $o_{t}$ for the exploration and question answering is not enough. Therefore, the agent should take advantage of its previous partial observations to form memory.
@@
To maximize its rewards, the exploration policy should guide the agent to spend more time in “informative” rooms, where the likelihood of encountering valuable information is higher. This is crucial because objects do not spend equal amounts of time in each room, meaning that not all sub-graphs are equally informative. Therefore, the agent needs to utilize its long-term memory to maintain and retrieve information about the maze layout and object locations, which is essential for both exploration and answering questions.
@@
In this work, “question answering” is approached as a form of KG reasoning, specifically restricted to answering one-hop queries. These queries involve predicting the tail in a given RDF triple $(s,r,?)$, where $s$ is the subject, $r$ is the relation, and the goal is to predict the unknown tail. This type of reasoning involves traversing the graph from a known entity using a specified relation.
@@
When the environment randomly samples objects (RDF triples) to generate questions, not all objects have an equal probability of being sampled. This bias in sampling affects both the exploration policy and the question-answering function. The agent must account for this sampling bias by focusing its exploration on areas with a higher likelihood of being queried.
@@
The Rooms environment is highly configurable, allowing for a wide range of experimental setups. Researchers can adjust various parameters such as the frequency of questions asked, the total number of questions per episode, the number of rooms, the layout of the rooms, the number and types of objects, and the number of steps per episode, etc. This flexibility enables the creation of diverse scenarios and difficulty levels, facilitating comprehensive studies on the effectiveness of different strategies.
@@ 3.2 Agents and Their Learning Objectives
Baseline Agent
The baseline agent utilizes the history of the past $\tau$ observations, denoted as $\bm{h}^{(\tau)}_{t}=(o_{t-\tau+1},\ldots,o_{t})$, to represent the current estimated state $\hat{s}_{t}$. This approach is the most straightforward method for handling a POMDP. For the question-answering function $f^{qa}$, the agent employs a simple strategy: using the most recent relevant memory to answer the question. This strategy assumes that the most recent observation contains the most accurate and relevant information for responding to queries about the environment.
In this setup, solving the POMDP involves optimizing the exploration policy $\pi^{\text{explore}}$, where the reward function the question-answering function $f^{qa}$. Thus, the “current state” for $\pi^{\text{explore}}$ is $\bm{h}^{(\tau)}_{t}$, and the “next state” is the updated $\bm{h}^{(\tau)}_{t+1}$, where the oldest observation is removed, and the new observation is added to the sequence.
HumemAI Agent
The agent with human-like memory systems (hereinafter HumemAI) introduces a function, $f^{\text{mm}}$, known as the memory management function. This agent features a memory system $\bm{M}_{t}$, which includes both short-term and long-term memory components, $\bm{M}^{\text{short}}_{t}$ and $\bm{M}^{\text{long}}_{t}$, respectively. The long-term memory system is further divided into episodic ($\bm{M}^{\text{episodic}}_{t}$) and semantic ($\bm{M}^{\text{semantic}}_{t}$) memory.
The function $f^{\text{mm}}$ first transfers the current partial observation to the short-term memory system by adding the relation qualifier current_time along with its value. For example, (Agent, atLocation, Home Office) at $t=42$ becomes (Agent, atLocation, Home Office, {current_time: 42}). The function then determines whether this short-term memory should be (1) moved to the episodic part of the long-term memory, (2) moved to the semantic part of the long-term memory, or (3) forgotten. If option (1) is selected, the short-term memory is stored in the episodic memory as (Agent, atLocation, Home Office, {timestamp: 42}). If option (2) is chosen, it is stored as (Agent, atLocation, Home Office, {strength: 1}).
@@
[…]
We aim to learn the function $f^{\text{mm}}$ using RL. To achieve this, we set the exploration policy $\pi^{\text{explore}}$ using a handcrafted function $f^{\text{explore}}$, which directs the agent to “go to random rooms while avoiding walls, utilizing the long-term memory.” For the question-answering function $f^{qa}$, we use the strategy of “looking up the most recent relevant or the strongest relevant memory.” By doing this, we have repurposed the original learning objective to learning the memory management policy $\pi^{\text{mm}}(a^{\text{mm}}_{t}|\bm{M}^{\text{long}}_{t-1},\bm{M}^{\text{short}}_{t})$, where $a^{\text{mm}}_{t}$ represents one of the three memory management actions. As for $\pi^{\text{mm}}$, the “current state” is the current memory and the “next state” is the next memory. Figure 2 shows the four steps involved in training.
@@
The learning process involves two distinct phases. First, we focus on learning $\pi^{\text{mm}}(a^{\text{mm}}_{t}|\bm{M}^{\text{long}}_{t-1},\bm{M}^{\text{short}}_{t})$. This phase ensures that the agent can effectively manage its long-term memory by deciding which observations to store as episodic or semantic memories, or which to forget. Once the memory management function is optimized, we then proceed to the second phase, which involves learning the exploration policy $\pi^{\text{explore}}(a^{\text{explore}}_{t}|\bm{M}^{\text{long}}_{t})$, by taking advantage of the function $f^{\text{mm}}$ learned in the first phase. In this phase, the agent uses its long-term memory to make informed navigation decisions within the environment. Here, the “current state” for $\pi^{\text{explore}}$ is the current long-term memory $\bm{M}^{\text{long}}_{t}$, and the “next state” is $\bm{M}^{\text{long}}_{t+1}$, updated by $\bm{M}^{\text{long}}_{t+1}=f^{\text{mm}}(\bm{M}^{\text{long}}_{t},\bm{M}^{\text{short}}_{t+1},a^{\text{mm}}_{t+1})$
An advantage of this two-phase approach is that the human-like long-term memory can effectively capture the most likely snapshot of the environment (the most likely belief state). This enables the agent to make well-informed decisions on exploration and question-answering.

### v2

@@ Abstract
Humans observe only part of their environment at any moment but can still make complex, long-term decisions thanks to our long-term memory. To test how an AI can learn and utilize its long-term memory, we have developed a partially observable Markov decision processes (POMDP) environment, where the agent has to answer questions while navigating a maze. The environment is completely knowledge graph (KG) based, where the hidden states are dynamic KGs. A KG is both human- and machine-readable, making it easy to see what the agents remember and forget. We train and compare agents with different memory systems, to shed light on how human brains work when it comes to managing its own memory. By repurposing the given learning objective as learning a memory management policy, we were able to capture the most likely hidden state, which is not only interpretable but also reusable. The code is open-sourced at https://github.com/humemai/agent-room-env-v2-lstm.
@@ 1 Introduction
The cognitive science theory suggests that humans have long-term memory (Tulving 1983, 1985; Tulving and Thomson 1973). This allows humans to solve various tasks, although they can only observe the world partially. For example, when trying to navigate through a familiar city, humans rely on stored memories of the layout and landmarks. This enables the recall of past experiences, aiding in decision-making and problem-solving even when immediate sensory input is limited. Similarly, when answering a question, humans draw on long-term memory to retrieve relevant facts, concepts, and previous knowledge.
Although there have been many studies, we still lack a comprehensive understanding of how our long-term memory is managed. Research has identified mechanisms for encoding, storing, and retrieving memories, but the precise regulation and optimization of these processes remain unclear (Artuso and Palladino 2019; Squire 2009).
@@
  • Optimal Memory Management Policy: We show that the best agent is the one that repurposes the given learning objectives as learning a memory management policy, allowing itself to capture the most likely hidden state, which is not only interpretable but also reusable.
@@ 2.1 A Machine With Human-Like Memory Systems
Recent advances in cognitive science and AI have inspired the development of machines with memory systems inspired by human memory. Kim et al. (Kim et al. 2022, 2023) introduced an agent model that explicitly incorporates both semantic and episodic memory, demonstrating that such an agent outperforms those with only one type of long-term memory. Their environment is designed to challenge the agent’s ability to encode, store, and retrieve memories effectively.
@@
To computationally model memory, Kim et al. (Kim et al. 2022, 2023) add relation qualifier key-value pairs to RDF (Resource Description Framework) triples to make them RDF quadruples (Lanthaler, Wood, and Cyganiak 2014; Hartig et al. 2024). For example, given a memory triple (Agent, atLocation, Library), they would turn this into something like (Agent, atLocation, Library, {timestamp: 42}), and (Agent, atLocation, Library, {strength: 2}), to turn them into episodic and semantic memories, respectively.
@@ 2.2 Reinforcement Learning (RL) in Partially Observable Markov Decision Process (POMDP)
RL is useful when supervision of every action is not feasible, such as in memory management. Solving an RL problem is often formalized as a Markov Decision Process (MDP) or a Partially Observable MDP (POMDP), which account for partial observability.
@@
Taking advantage of belief states can help solve POMDP problems. The environment starts in an initial state distribution. At time step $t$, the agent maintains a belief state $\mathrm{b}_{t}$, which is a probability distribution over the state space $\mathcal{S}$. The agent then chooses an action $\mathrm{a}_{t}\in\mathcal{A}$ based on its current belief state $\mathrm{b}_{t}$ according to a policy $\pi:\mathcal{B}\rightarrow\mathbb{P}(\mathcal{A})$, where $\mathcal{B}$ denotes the belief space. After taking action $\mathrm{a}_{t}\sim\pi(\mathrm{a}_{t}\mid\mathrm{b}_{t})$, the agent observes an observation $\mathrm{o}_{t+1}\in\Omega$. Using the observation $\mathrm{o}_{t+1}$ and the action $\mathrm{a}_{t}$, the agent updates its belief state to $\mathrm{b}_{t+1}$.
@@
where $r_{t}=R(\mathrm{s}_{t},\mathrm{a}_{t})$, $\mathrm{a}_{t}\sim\pi(\mathrm{a}_{t}\mid b_{t})$, $\mathrm{s}_{t+1}\sim T(\mathrm{s}_{t},\mathrm{a}_{t})$, and $\mathrm{o}_{t+1}\sim O(\mathrm{o}_{t+1}\mid\mathrm{s}_{t+1},\mathrm{a}_{t})$. The belief state $b_{t+1}(\mathrm{s}_{t+1})$ is updated using the current belief state $\mathrm{b}_{t}$, the action $\mathrm{a}_{t}$, and the observation $\mathrm{o}_{t+1}$, along with the observation model $O$ and the transition model $T$. The update is done by $b^{\prime}(s^{\prime})=\eta\,O(o^{\prime}\mid s^{\prime},a)\sum_{s}T(s^{\prime}\mid s,a)b(s)$, where $O(o^{\prime}\mid s^{\prime},a)$ is the observation likelihood, $T(s^{\prime}\mid s,a)$ is the transition model, and $\eta$ is a normalizing constant.
Updating the belief state is computationally expensive because it involves maintaining a probability distribution over a potentially vast state space, requiring summing over all possible states while applying transition and observation probabilities. This process is further complicated by the need for the world model, i.e., the observation model and transition function.
A simple way to mitigate this computational expense, especially in the era of deep learning, is to use the history of the past $\tau$ observations $\bm{h}^{(\tau)}_{t}=(o_{t-\tau+1},\ldots,o_{t})$ as the estimated state $\hat{s}_{t}$, with which we learn the policy $\pi(\mathrm{a}_{t}\mid\mathrm{\hat{s}}_{t})$. Although not perfect, this alleviates the need to maintain or update the belief state $\mathrm{b}_{t}$.
@@ 3.1 The Rooms Environment
The Rooms Environment provides a simulation for evaluating AI agents with long-term memory. Unlike the previous work, this environment consists of multiple interconnected rooms, each containing various types of objects. Static objects, such as beds and tables, remain fixed in their locations and do not move. Independent objects, like Bob and Alice, move between rooms according to probabilistic rules that dictate their movements. Dependent objects, which cannot move by themselves, may be transported by independent objects with a certain probability that depends on the current state of the room. Finally, the agent represents the RL agent that navigates between rooms based on its exploration policy. This configuration allows for the study of how agents manage and utilize long-term memory to make decisions in a dynamic and partially observable environment.
In this environment, hidden states ($s_{t}$) are represented as dynamic KGs that capture the entire rooms and the locations of all objects within them. Each room and its contents form a sub-graph of this overall KG. Since the agent can only be present in one room at a time, it can only observe the objects within that particular room. This partial observability requires the agent to utilize its long-term memory.
@@
Figure 1 and 2 show an example hidden state $s_{t}$ and partial observation $o_{t}$. In this example, there are 32 rooms, i.e., Office room, Kitchen, …, and 25 objects, i.e., Coat, Towel, …, Agent. Therefore, the number of all hidden states is $|S|=N_{\text{rooms}}^{(N_{\text{objects}}-N_{\text{static objects}})}=32^{25-8}\approx 3.8\times 10^{25}$, since every object is located in one of $N_{\text{rooms}}$ and the static objects do not move. The number of possible sub-graphs, i.e., partial observation, that the agent can observe is $2^{(N_{\text{objects}}-N_{\text{static objects}})}\times N_{\text{rooms}}=2^{25-4}\approx 6.7\times 10^{7}$, since the agent can either see or do not see an object in the room. Although the environment might look simple, we have a huge state space, which can be too much for typical table-based RL methods.
@@
In this example, the partial observation (sub-graph of the hidden dynamic KG) is { (Closet, north, Nursery), (Closet, east, Wall), (Closet, south, Wall), (Closet, west, Sauna), (Agent, atLocation, Closet), (Sam, atLocation, Closet) }. The hidden state is a much bigger KG that contains this sub-graph.
@@
Figure 1: An illustration of a hidden state $s_{t}$ (in white) and partial observation $o_{t}$ (in gray).
@@
Figure 2: A hidden state $s_{t}$ (in white) and partial observation $o_{t}$ (in gray) represented as a KG.
@@
The exploration policy, $\pi^{\text{explore}}(a^{\text{explore}}|o_{t})$, allows the agent to navigate through the room by taking one of five possible actions: North, East, South, West, or Stay. As the agent explores, the environment intermittently poses questions about the object locations. Consequently, the agent must not only learn an effective exploration policy, but also develop a robust question-answering function, $answer=f^{\text{qa}}(o_{t},question)$. The question-answering function also serves as the reward function: the agent receives a reward of +1 for each correctly answered question and 0 otherwise. Although straightforward, using $o_{t}$ for the exploration and question answering is not enough. Therefore, the agent should take advantage of its previous partial observations to form memory.
@@
To maximize its rewards, $\pi^{\text{explore}}$ should guide the agent to spend more time in “informative” rooms, where the likelihood of encountering valuable information is higher. This is crucial because objects do not spend equal amounts of time in each room, meaning that not all sub-graphs are equally informative. Therefore, the agent needs to utilize its long-term memory to maintain and retrieve information about the maze layout and object locations, which is essential for both exploration and answering questions.
@@
In this work, “question answering” is approached as a form of KG reasoning, specifically restricted to answering one-hop queries. These queries involve predicting the tail in a given RDF triple $(s,r,?)$, where $s$ is the subject, $r$ is the relation, and ???? is the unknown tail. This type of reasoning involves traversing the graph from a known entity using a specified relation.
@@
When the environment randomly samples objects to generate questions, not all objects have an equal probability of being sampled. This bias in sampling affects both the exploration policy and the question-answering function. The agent must account for this sampling bias by focusing its exploration on areas with a higher likelihood of being queried.
@@
The Rooms environment is highly configurable, allowing for a wide range of experimental setups. Researchers can adjust various parameters such as the question frequency, the total number of questions, the number of rooms, the room layout, the number and types of objects, and the number of steps per episode, etc. This flexibility enables the creation of diverse scenarios and difficulty levels.
@@ 3.2 Agents and Their Learning Objectives
Baseline
The baseline agent utilizes the history of the past $\tau$ observations, denoted as $\bm{h}^{(\tau)}_{t}=(o_{t-\tau+1},\ldots,o_{t})$, to represent the current estimated state $\hat{s}_{t}$. This approach is the most straightforward method. For the question-answering function $f^{\text{qa}}$, the agent employs a simple strategy: use the most recent relevant memory to answer the question. This strategy assumes that the most recent observation contains the most accurate and relevant information for the question.
In this setup, solving the POMDP involves optimizing the exploration policy $\pi^{\text{explore}}(\mathrm{a}^{\text{explore}}_{t}\mid\bm{h}^{(\tau)}_{t})$. Thus, the “current state” for $\pi^{\text{explore}}$ is $\bm{h}^{(\tau)}_{t}$, and the “next state” is the updated $\bm{h}^{(\tau)}_{t+1}$, where the oldest observation is removed, and the new observation is added to the sequence.
HumemAI
The agent with human-like memory systems (hereinafter HumemAI) introduces a function, $f^{\text{mm}}$, “the memory management function”. This agent has a memory system $\bm{M}_{t}$, which includes both short-term and long-term memory components, $\bm{M}^{\text{short}}_{t}$ and $\bm{M}^{\text{long}}_{t}$, respectively. The long-term memory is further divided into episodic ($\bm{M}^{\text{episodic}}_{t}$) and semantic ($\bm{M}^{\text{semantic}}_{t}$).
[…]
@@
In essence, the long-term memory system is updated at each step according to the equation $\bm{M}^{\text{long}}_{t}=f^{\text{mm}}(\bm{M}^{\text{long}}_{t-1},\bm{M}^{\text{short}}_{t},a^{\text{mm}}_{t})$. The agent begins with an empty long-term memory and progressively updates it with each new observation.
We aim to learn the function $f^{\text{mm}}$ using RL. To achieve this, we “fix” the exploration policy $\pi^{\text{explore}}$ with a handcrafted function $f^{\text{explore}}$, which directs the agent to “go to random rooms while avoiding walls, utilizing the long-term memory.” For the question-answering function $f^{\text{qa}}$, we use the strategy of “look up the most recent or the strongest relevant memory.” By doing this, we have repurposed the original learning objective into learning the memory management policy $\pi^{\text{mm}}(a^{\text{mm}}_{t}|\bm{M}^{\text{long}}_{t-1},\bm{M}^{\text{short}}_{t})$, where $a^{\text{mm}}_{t}$ represents one of the three memory management actions. As for $\pi^{\text{mm}}$, the “current state” is the current memory and the “next state” is the next memory. Figure 3 shows the four steps involved in training.
@@
The learning process involves two distinct phases. First, we focus on learning $\pi^{\text{mm}}(a^{\text{mm}}_{t}|\bm{M}^{\text{long}}_{t-1},\bm{M}^{\text{short}}_{t})$. This phase ensures that the agent can effectively manage its long-term memory by deciding which observations to store as episodic or semantic memories, or which to forget. Once this is optimized, we proceed to the second phase, which involves learning the exploration policy $\pi^{\text{explore}}(a^{\text{explore}}_{t}|\bm{M}^{\text{long}}_{t})$, by taking advantage of the function $f^{\text{mm}}$ learned in the first phase. In this phase, the agent uses its long-term memory to make informed navigation decisions within the environment. Here, the “current state” for $\pi^{\text{explore}}$ is the current long-term memory $\bm{M}^{\text{long}}_{t}$, and the “next state” is $\bm{M}^{\text{long}}_{t+1}$, updated by $\bm{M}^{\text{long}}_{t+1}=f^{\text{mm}}(\bm{M}^{\text{long}}_{t},\bm{M}^{\text{short}}_{t+1},a^{\text{mm}}_{t+1})$
An advantage of this two-phase approach is that the human-like long-term memory can effectively capture the most likely snapshot of the environment (the most likely hidden state). This enables the agent to make well-informed decisions on exploration and question-answering.